OpenStudy (anonymous):

sin 3Q + sin Q = 2 sin 2Q cos Q

OpenStudy (anonymous):

Are we ask to prove it ???

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Any guess ????

OpenStudy (anonymous):

What do u mean by any guess?

OpenStudy (anonymous):

i know i can split sin 3Q= sin(2Q+Q)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

but i am stucked after that.

OpenStudy (anonymous):

just a moment

OpenStudy (anonymous):

Here at right hand side just use trigonometric formula sin(a) + sin(b) = 2* sin((a+b)/2) * cos((a-b)/2 Hope this will help you

OpenStudy (anonymous):

i dont see any relation?

OpenStudy (anonymous):

We are having left hand side as sin3Q + sinQ compare it with sina + sinb So a = --- ? b = --- ?

OpenStudy (anonymous):

so a = 3Q & b = Q plug in that formula .,,& simplify it

OpenStudy (anonymous):

can someone please show me the steps? sin3QcosQ + cos3QsinQ ?

OpenStudy (anonymous):

Q=x sin(3x)+sin(x)=2sin(2x)cos(x) sin(3x)+sin(x)=2sin((3x+x)/2)cos((3x-x)/2)= =2sin(2x)cos(x)!! I use the trigonometric (sin(a+b)=2sin((a+b)/2)cos((a-b)/2)

OpenStudy (anonymous):

SIN(2Q+Q)+SIN_(2Q-Q)

OpenStudy (anonymous):

phaniraj61: your answer is 2sinQcosQ ?

OpenStudy (whpalmer4):

You're trying to show that \[\sin(3x) + \sin(x) = 2\cos(x)\sin(2x)\] You can expand \(\sin(3x)\) via identity \(\sin(u+v) = \sin(u)\cos(v) + \cos(u)\sin(v)\) with \(u=2x,v=x\): \[\sin(3x) = \sin(2x+x) = \sin(2x)\cos(x) + \cos(2x)\sin(x)\]our left side becomes \[\sin(2x)\cos(x) + \cos(2x)\sin(x)+\sin(x) = 2\cos(x)\sin(2x)\]Next, use the identity \[\cos(2x) = 2\cos^2(x)-1\] \[\sin(2x)\cos(x) + (2\cos^2(x)-1)\sin(x)+\sin(x) = 2\cos(x)\sin(2x)\]\[\sin(2x)\cos(x) + 2\cos^2(x)\sin(x)\cancel{ -\sin(x)+\sin(x) }= 2\cos(x)\sin(2x)\]Now just use \[\sin(2x) = 2\sin(x)\cos(x)\]to convert the middle term into the right form and you're all done.